OpenStudy (anonymous):

The area of a rectangle that is three times as long as it is wide if w is the width of the rectangle.

OpenStudy (shadowlegendx):

Do you have any optional choices?

OpenStudy (anonymous):

i don't have. i guess it is an equation?

OpenStudy (shadowlegendx):

A visual example is best for solving this type of question

OpenStudy (shadowlegendx):

Remember that \[A = wl\]

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

it is an identification type.

OpenStudy (ahsome):

Do you know the area in the first place?

OpenStudy (shadowlegendx):

\[A = w(w^{3} )\]

OpenStudy (anonymous):

nope it is not given.

OpenStudy (unklerhaukus):

make sure to substitute correctly, A = WL ( L = 3W ) = W(3W) ( now simplify the expression ) =